OpenStudy (hellomynameiskam):

What is the ratio 15 : 75 in lowest terms? A. 1 : 5 B. 2 : 5 C. 1 : 4 D. 3 : 5

OpenStudy (hellomynameiskam):

@Godlovesme

OpenStudy (godlovesme):

have u tried like i showed u earlier? :)

OpenStudy (hellomynameiskam):

yea. im horrible at writting so i cant really read my writing

OpenStudy (mathteacher1729):

This problem is asking you to reduce the fraction \[\frac{15}{75}. \] Do you know how to reduce a fraction like this?

OpenStudy (hellomynameiskam):

no. im kinda new at reproducing fractions.

OpenStudy (anonymous):

1/5 is the answer

OpenStudy (mathteacher1729):

Basically you factor the top and bottom of the fraction and cross off any numbers in common: Example: \[\frac{5}{10}=\frac{5}{5\cdot2}=\frac{1}{2}\]